• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Calarco

    Log entry time 16:03:05 on June 2,2000

    Entry number 42884

    keyword=Shift Summary

    Continuation of running at Theta_CM = 135, Phi_CM = 180
    .
    Run 1683 ... 0.9M events
    .
    Pointing runs:
    Run 1684 ... 0.4M events, 10 cm dummy
    Run 1685 ... 0.2M events, 15 cm dummy
    Run 1686 ... 0.2M events, 4 cm dummy
    Run 1687 ... 0.1M events, carbon
    Run 1688 ... harp scan, 923.data
    .
    Asked for access at 10:00 to work on trigger, finished at 13:30. We think we
    have solved the problem of double peaking in the self-timing TDC spectra
    (see entry by Bob Michaels). I have asked MCC to close the hall and called
    Eugene for a Moeller measurement. During the access we also changed the
    HRSh angle to 34.73 deg and the momentum to 0.7418 GeV for the kinematics
    at Theta_CM = 155, Phi_CM = 0.
    .
    Runs 1689 - 1692 are timing tests
    .
    Run 1693 ... cosmics to check self-timing
    Run 1694 ... more cosmics, all PS set to 1
    .

    We have not had beam since 10:00 due to accelerator problems. Finally, at
    15:40, we have beam again. Eugene will do his Moeller measurements.

    I'm outta here!!